an equilateral triangle has a perimeter in meters of 6a + 9. what is an expression in simplest form for the length of each side in meters
Answer by rothauserc(4718)

An equilateral triangle has all three sides(s) equal, therefore perimeter(P) is the following formula,
P = 3*s
We are given that P = 6*a + 9, therefore
3*s = 6*a + 9
Divide both sides of = by 3
s = 2*a + 3 meters